The perimeter of a rectangle is 
120.
The length is 5 more then 4 
times the width.
Find the length and the width.

Answer by stanbon(75887)

The perimeter of a rectangle is
120.
The length is 5 more then 4
times the width.
Find the length and the width.
-----
Let the width be "x":
Then length = 4x+5
----
Equation:
P = 2(L + W)
120 = 2(4x+5 + x)
60 = 5x + 5
5x = 55
x = 11 (width)
-----
4x+5 = 49 (length)
======================
Cheers,
Stan H.
==================